Circular No. 7903

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

1ES 1959+650 T. Weekes, Harvard-Smithsonian Center for Astrophysics (CfA), com-

municates: “C. Dowdall and P. Moriarty, Galway-Mayo Institute of Technology; and K. Kosack, Washington University, report (for the VERITAS collaboration) a strong detection in TeV γ-rays (peak energy > 400 GeV) of the x-ray-selected BL-Lac object 1ES 1959+650 at a redshift z = 0.048. The observations were made with the VERITAS 10-m γ-ray telescope at the Whipple Observatory on Mt. Hopkins on May 16.444–16.462 and 17.290– 17.457 UT. The detection is significant at the 13σ level. The γ-ray rate doubled over the two nights of observation; by the end of the observation period, it had reached a maximum of 2.5 times the Crab Nebula flux. A preliminary analysis of the energy spectrum indicates a hard spectrum with differential index −2.40 ± 0.14 from 260 GeV to 5.6 TeV. Observations at other wavelengths are encouraged; VERITAS observations will continue through May 21.”

SUPERNOVA 2002cx IN ANONYMOUS GALAXY T. Matheson, S. Jha, P. Challis, and R. Kirshner, CfA, report that a

spectrum (range 370–750 nm) of SN 2002cx (cf. IAUC 7902), obtained by P. Berlind on May 17.22 UT with the Mt. Hopkins 1.5-m telescope, shows it to be a peculiar type-Ia supernova before maximum light. The Si II (635.5 nm) and Ca II (395.1 nm) features are extremely weak or absent, but the Fe III lines at 430 and 500 nm are present. The spectrum is almost identical to that of SN 1991T at 7 days before maximum (see Filippenko et al. 1992, Ap.J. 384, L15).

SUPERNOVA 2002cv IN NGC 3190 W. Li, University of California at Berkeley, reports that no new object

appears at the location of SN 2002cv (cf. IAUC 7901) on KAIT images taken in bands B, V, and R on May 9, 13, or 14, in B and V on May 15, and in I on May 9. An object is, however, visible in an I-band image taken in poor seeing with the Lick Observatory 1.0-m Nickel telescope on May 14 (I ∼ 18.5), and SN 2002cv is present at R = 20.3 and I = 17.2 (photometry from Nickel-telescope images) on May 15.23 UT, located at α = 10^{h}18^{m}03^{s}.68, δ = +21^{o}50′06′′.0 (equinox 2000.0; astrometry from a KAIT I-band image), which is 27′′.1 west and 10′′.5 north of the nucleus of NGC 3190. The new object is superimposed on the dust lane of NGC 3190.

2002 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green